An empty plane whose mass is 30, 400Kg has a maximum takeoff acceleration of 1.20m/s2. What is its maximum acceleration when it is carrying a load of 8200Kg? Ignore friction.
1
Expert's answer
2021-09-21T06:33:11-0400

The force (thrust) of the engines, calculated for curb weight:


"F=ma=30400·1.2=36480\text{ N}."

The acceleration with the load is


"a_L=\frac F{m+M}=\frac{36480}{30400+8200}=0.95\text{ m/s}^2."